OpenStudy (anonymous):

Solve for the variable. a + 5.1 = 10.9 A. 5.8 B. -5.8 C. 16 D. -16

OpenStudy (anonymous):

a=5.8, or A

OpenStudy (hitaro9):

The best way to think of these problems is to say "What plus 5.1 =10.9" You can try a few different of the options. will 16 + 5.1 = 10.9? No, so that can't be it. If you go through the options, you'll see A, 5.8, is the only one that fits.

OpenStudy (hitaro9):

Is there anything you're having trouble with in specific with these problems? You can think of the a as a question mark ? if it makes it easier.

OpenStudy (nikato):

Or because you see that the problem is +5.1 and you're trying to get a isolated on one side. To do this you must undo the +, so you must -. So a+5.1=10.9 -5.1. -5.1 ---------------- a=? , which is whatever 10.9-5.1 is